The molecular weight of calcium carbonate CaCO3

 is:

(A) 80

(B) 90

(C) 100

(D) 110

 

Grade:12th pass

3 Answers

Vikas TU
3 years ago
Dear student 
Molecular weight of Ca = 40
Molecular weight of C = 12 
Molecular weight of O = 16 
Molrcular weight of CaCO 3 = 100 g/mol
 
Vikas Amritiya
3 years ago

Molecular weight is the mass of whole molecule.

Steps:- (1) Chemical formula of calcium carbonate CaCO3
(2) Mass of Calcium40
     Mass of Carbon12
     Mass of Oxygen16
(3) Multiply mass of each element by number of atoms of that element.
(4) Add all the masses obtained after multiplication.
CaCO3
=1×40+1×12+3×16
=40+12+48
=100.
irshad hussain
2 years ago
Correct option is
C

100

Molecular weight is the mass of whole molecule.

Steps:- (1) Chemical formula of calcium carbonate CaCO3
(2) Mass of Calcium40
     Mass of Carbon12
     Mass of Oxygen16
(3) Multiply mass of each element by number of atoms of that element.
(4) Add all the masses obtained after multiplication.
CaCO3
=1×40+1×12+3×16
=40+12+48
=100.